OpenStudy (anonymous):

Evaluate the limit. lim x---> 0 sqrt(x+1) - sqrt(2x+1)/sqrt(3x+4) - sqrt(2x+4) I got lim x--> 0 -sqrt(3x+4)-sqrt(2x+4)/sqrt(x+1)+sqrt(2x+1) Is there anyway to simplify it down further?

OpenStudy (anonymous):

\[\lim_{x\to0}\frac{\sqrt{x+1}-\sqrt{2x+1}}{\sqrt{3x+4}-\sqrt{2x+4}}\cdot\frac{\sqrt{x+1}+\sqrt{2x+1}}{\sqrt{x+1}+\sqrt{2x+1}}\cdot\frac{\sqrt{3x+4}+\sqrt{2x+4}}{\sqrt{3x+4}+\sqrt{2x+4}}\\ \lim_{x\to0}\frac{x+1-(2x+1)}{3x+4-(2x+4)}\cdot\frac{\sqrt{3x+4}+\sqrt{2x+4}}{\sqrt{x+1}+\sqrt{2x+1}}\\ \lim_{x\to0}-\frac{x}{x}\cdot\frac{\sqrt{3x+4}+\sqrt{2x+4}}{\sqrt{x+1}+\sqrt{2x+1}}\\ \vdots \]

OpenStudy (anonymous):

Now you can just plug in the 0.

OpenStudy (anonymous):

Ok, thanks

OpenStudy (anonymous):

yw